Let X denote the amount of time (in minutes) a postal clerk spends with his/her customer. The time is known to have an exponential distribution with an average amount of time equal to 4 minutes.

i)Find the probability that a clerk spends four to five minutes with a randomly selected customer.

ii)Half of all the customers are finished within how long? (Find the median)

iii)Which is larger, the mean or the median?


1
Expert's answer
2021-08-10T15:21:48-0400

"X" is a continuous random variable since time is measured. It is given that "\mu=4" minutes. 

"X\sim Exp(\lambda)"


"\lambda=\dfrac{1}{\mu}=\dfrac{1}{4}=0.25"

i)

"P(4<X<5)=P(X<5)-P(X<4)"

"=1-e^{-\lambda(5)}-(1-e^{-\lambda(4)})"

"=e^{-0.25(4)}-e^{-0.25(5)}=e^{-1}-e^{-1.25}"

"\approx0.081375"

ii)


"P(X<T)=0.5"

"1-e^{-0.25(T)}=0.5"

"e^{-0.25(T)}=0.5"

"0.25(T)=\ln 2"

"T=4\ln 2"

"T\approx2.77\ minutes"

iii) The theoretical mean is  "4" minutes. We have just calculated the median (or 50th percentile) as  "2.77" minutes. So the mean is larger.